I can't wrap my brain around this word problem.
Layoffs. 10% of the employees in the aerospace division will be laid off and 15% of the employees in the agriculture division will be laid off. Altogether 12% of 3000 employees in these two divisions will be laid off. How many employees are in each division?
__________________
12% of 3000= 360
10%=10/100=.10
15%=15/100=.15

Answer by stanbon(75887)

I can't wrap my brain around this word problem.
Layoffs. 10% of the employees in the aerospace division will be laid off and 15% of the employees in the agriculture division will be laid off. Altogether 12% of 3000 employees in these two divisions will be laid off. How many employees are in each division?
----------------
Quantity Equation: s + a = 3000
Percentage Equatio: 0.10s + 0.15a = 0.12*3000
---------------------------------
Rearrange the equations for elimination:
10s + 10a = 30000
10s + 15a = 12*3000
----------------
Subtract the 1st from the 2nd to get:
5a = 6000
a = 1200 (# of agriculture division layoffs)
s = 1800 (# of aerospace division layoffs)
======================
Cheers,
Stan H.